OpenStudy (anonymous):

how to solve 49x^2+25

OpenStudy (anonymous):

49x^2+25=0 49x^2=-25 x^2=-25/49 x=sqrt(-25/49)

jhonyy9 (jhonyy9):

x_1,2= +/- (5i)/7